1. Roronoa Zoro runs forward, from rest, with an acceleration of 8 m/s2.  If he covered 25 m, what was his final velocity?




  1. A rollerskater experiences an acceleration of 2.2 m/s2. As a result, she accelerates from rest to 12 m/s. How much distance did she travel during that acceleration?






  1. A hat is dropped from a 44 m tall building. How fast will it be going when it strikes the ground? (Acceleration is -9.8 m/s2)






  1. The Lizard chases after Spider-Man and experiences an acceleration of 5.2 m/s2. As a result, he accelerates from rest to 14.8 m/s. How much distance did the Lizard cover?
1
Expert's answer
2022-01-05T04:21:16-0500

1)

"v^2=v_0^2+2ad,""v=\sqrt{2ad}=\sqrt{2\times8\ \dfrac{m}{s^2}\times25\ m}=20\ \dfrac{m}{s}."

2)

"v^2=v_0^2+2ad,""d=\dfrac{v^2-v_0^2}{2a}=\dfrac{(12\ \dfrac{m}{s})^2}{2\times2.2\ \dfrac{m}{s^2}}=32.7\ m."

3)

"v^2=v_0^2+2gh,""v=\sqrt{2gh}=\sqrt{2\times(-9.8\ \dfrac{m}{s^2})\times(-44\ m)}=29.4\ \dfrac{m}{s}."

4)

"d=\dfrac{v^2-v_0^2}{2a}=\dfrac{(14.8\ \dfrac{m}{s})^2}{2\times5.2\ \dfrac{m}{s^2}}=21.1\ m."